OpenStudy (anonymous):

How would I write the following problem in a Linear equation? A mom bought 28 packages of Reese’s Cups with 88 total cups. Each king size package holds 4 cups. Each regular size package holds two cups. How many of each size did she buy?

OpenStudy (anonymous):

k+r=28 4k+2r=88

OpenStudy (amistre64):

(14+n) + (14-n) = 28 4(14+n) + 2(14-n) = 88 solve for n